OpenStudy (matttttt):

Which of the following measurements is equal to 0.373 kL? A)373000 L B)3730 L C)373 L D)37.3 L

OpenStudy (lfreeman1285):

Because all of the answers are in liters, the best course of action would be to convert 0.373 kiloliters (kL) into liters (L), as shown below. \[\frac{ 0.373 kL }{ 1 } \times \frac{ 1000 L }{ 1 kL } = 373 L\] There are 1000 L in a kL, so multiplying 0.373 kL by 1000 L / kL cancels out the kL, leaving liters as the unit of measure. Therefore, the answer is C) 373 L

OpenStudy (geobit):

yeah it is.